OpenStudy (anonymous):

find the derivative \[s=t^7-\cos t-7t \sin t-7 \cos t\]

OpenStudy (anonymous):

is it just \[s=7t^6 \sin t -7t \cos t -7 \sin t \]

OpenStudy (anonymous):

\[s=7t ^{6}+\sin t - (7t \times \cos t + 7 \times \sin t) + 7\sin t\] \[7t ^{6} + \sin t -7t \cos t -7\sin t + 7\sin t\]